OpenStudy (anonymous):

Determine the value(s) for which the rational expression is undefined. -2y/7y-2

OpenStudy (anonymous):

you can't define it if 7y-2=0 (can't divide by 0, right?)

OpenStudy (anonymous):

so, it' s undefined if: 7y=2 or y=2/7